# Quantum Electrodynamics of graphene Landau levels in a deep-sub-wavelength hyperbolic phonon polariton cavity

### v1

@@ V.4 Two-transition model
Fig. 8(c) shows $-{\rm Im}~{}{\chi}_{\rm nn}(q_{\parallel},\omega)$ as a function of $\hbar v_{\rm F}/\ell_{B}$, for a small wavevector $\bar{q}_{\parallel}=0.05~{}{\rm nm}^{-1}$ and $L_{z}=50~{}{\rm nm}$. In this regime, where $q_{\parallel}L_{z}\gg 1$, only the Rabi frequency with $n_{z}=0$ remains significant, as the contributions from other $n_{z}$ modes rapidly decay to zero, as evident in Fig. 7(a). This behavior allows us to simplify the analysis by focusing exclusively on the interaction between a specific cavity mode $(q_{\parallel},\bar{n}_{z},\bar{s})$ and a single LL transition$(\bar{n},\bar{\lambda})\rightarrow(\bar{n}^{\prime},\bar{\lambda}^{\prime})$. Under this approximation, the dielectric function ${\rm Re}~{}\epsilon^{(2)}_{\rm RPA}(q_{\parallel},\omega)$ can be expressed as:
@@ VI Conclusions
Here, we investigated the polaritonic spectrum of LLs in graphene embedded in an HPP cavity footnote , including non-local effects. Our analysis demonstrated that the non-local response introduces spectral features absent in purely local theories, such as coupling to LL transitions that are dipole-forbidden, underscoring its crucial role in accurately describing light-matter interactions at the nanoscale. Notably, our nonlocal theory identifies the vacuum Rabi frequency is maximized when the light’s momentum, $q_{\parallel}$, is comparable to the inverse of the magnetic length, $q_{\parallel}\ell_{B}\sim 1$. This insight underscores the importance of considering nonlocal effects for achieving optimal light-matter coupling in such systems. Finally, we find that the strong coupling between LL transitions in graphene and cavity modes can be effectively engineered under high magnetic fields and low temperatures. Ref. andolina2024, established that sub-wavelength cavities are described by a scalar potential $\phi$ rather then a vector potential. Starting from this insight, we developed a new theoretical framework to study C-QED of sub-wavelength cavities, by quantizing the scalar potential $\hat{\phi}$, assessing quantum vacuum fluctuations, and providing a way to calculate vacuum Rabi frequencies, without resorting to phenomenological compression factors. Our theory provides a quantitative tool to directly assess light-matter interactions in the strong and ultra-strong coupling regimes, offering a predictive tool for cavity design.
@@
Our approach would open new avenues for engineering the electronic properties of materials embedded in cavities. For example, it could enable the design of superconductivity in moiré materials through tailored cavity structures BlochReview , the modulation of topological phenomena such as the quantum Hall effect appugliese2022 ; enkner2024 , or the enhancement and control of interaction effects, including fractional quantum Hall states enkner2024b ; bacciconi2023 or correlated states in twisted bilayer graphene Cao in a cavity. Indeed, our microscopic theory of sub-wavelength cavities can be used to investigate how cavities can be employed to enhance Ebbesen19 ; Sentef_ScienceAdvances_2018 or induce schlawin2019 superconductivity.
@@ VI.1 Acknowledgments
G.M.A acknowledges useful discussion with I. Torre, L. Orsini, J. Sueiro, G. Mazza, Z. Bacciconi, D. De Bernardis, G. Chiriacò, M. Michael, A. Rubio, M. Ruggenthaler and F. Schlawin. G.M.A acknowledge funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie (Grant Agreement No. 101146870 – COMPASS). G.M.A and M.S. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002955 – CONQUER). B.T. acknowledges funding from “la Caixa” Foundation (ID 100010434). R.R. and M.P. were supported by the MUR - Italian Ministry of University and Research under the “Research projects of relevant national interest - PRIN 2020” - Project No. 2020JLZ52N (“Light-matter interactions and the collective behavior of quantum 2D materials, q-LIMA”) and by the European Union under grant agreement No. 101131579 - Exqiral and No. 873028 - Hydrotronics. F.H.L.K. acknowledges support from the ERC TOPONANOP under grant agreement no. 726001, the Gordon and Betty Moore Foundation through Grant GBMF12212, and the Government of Spain (FIS2016-81044; PID2019-106875GB-100; Severo Ochoa CEX2019-000910-S [MCIN/ AEI/10.13039/501100011033], PCI2021-122020-2A, and PDC2022-133844-100 funded by MCIN/AEI/10.13039/501100011033). This work was also supported by the European Union NextGenerationEU/PRTR (PRTR-C17.I1) and EXQIRAL 101131579, Fundació Cellex, Fundació Mir-Puig, and Generalitat de Catalunya (CERCA, AGAUR, 2021 SGR 014431656). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Additionally, the research leading to these results received funding from the European Union’s Horizon 2020 under grant agreements no. 881603 (Graphene flagship Core3) and 820378 (Quantum flagship). This material is based upon work supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7047. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force.
@@ Appendix B The density–density response function for graphene in a magnetic field
Employing Eqs. (101,103) we can express the element matrix of the density operator $\hat{n}_{\bm{q}_{\parallel}}$ as:

### v2

@@ V.4 Two-transition model
Here, we identify a regime where a given LL transition is coupled only to a single cavity mode, allowing for an effective simplified description. Indeed, Fig. 9(c) shows $-{\rm Im}~{}{\chi}_{\rm nn}(q_{\parallel},\omega)$ as a function of $\hbar v_{\rm F}/\ell_{B}$, for a small wavevector $\bar{q}_{\parallel}=0.05~{}{\rm nm}^{-1}$ and $L_{z}=50~{}{\rm nm}$. In this regime, where $q_{\parallel}L_{z}\gg 1$, only the Rabi frequency with $n_{z}=0$ remains significant, as the contributions from other $n_{z}$ modes rapidly decay to zero, as evident in Fig. 8(a). This behavior allows us to simplify the analysis by focusing exclusively on the interaction between a specific cavity mode $(q_{\parallel},\bar{n}_{z},\bar{s})$ and a single LL transition$(\bar{n},\bar{\lambda})\rightarrow(\bar{n}^{\prime},\bar{\lambda}^{\prime})$. Under this approximation, the dielectric function ${\rm Re}~{}\epsilon^{(2)}_{\rm RPA}(q_{\parallel},\omega)$ can be expressed as:
@@ VI Conclusions
Here, we investigated the polaritonic spectrum of LLs in graphene embedded in an HPP cavity, including non-local effects beyond dipole approximation. Our analysis demonstrated that the non-local response introduces spectral features absent in purely local theories, such as coupling to LL transitions that are dipole-forbidden, underscoring its crucial role in accurately describing light-matter interactions at the nanoscale. Notably, our nonlocal theory identifies the vacuum Rabi frequency is maximized when the light’s momentum, $q_{\parallel}$, is comparable to the inverse of the magnetic length, $q_{\parallel}\ell_{B}\sim 1$. This insight underscores the importance of considering nonlocal effects for achieving optimal light-matter coupling in such systems. Finally, we find that the strong coupling between LL transitions in graphene and cavity modes can be effectively engineered under high magnetic fields and low temperatures. Ref. andolina2024, established that sub-wavelength cavities are described by a scalar potential $\phi$ solving the electrostatic Poisson equation rather than a vector potential. Starting from this insight, we developed a new theoretical framework to study C-QED of sub-wavelength cavities, by quantizing the scalar potential $\hat{\phi}$, assessing quantum vacuum fluctuations, and providing a way to calculate vacuum Rabi frequencies, without resorting to phenomenological compression factors. Our theory provides a quantitative tool to directly assess light-matter interactions in the strong and ultra-strong coupling regimes, offering a predictive tool for cavity design.
@@
Our approach paves the way for engineering the electronic properties of materials embedded in cavities. Given the numerous predictions in this field gao2020 ; schlawin2019 ; wang2019 ; nguyen2023 ; zezyulin2022 ; lenk2022 ; arwas2023 ; pavosevic2022 , it would be important to revisit their conclusion in light of our new theoretical framework for sub-wavelength cavities. For example, it could enable the design of superconductivity in moiré materials through tailored cavity structures BlochReview , the modulation of topological phenomena such as the quantum Hall effect appugliese2022 ; enkner2024 , or the enhancement and control of interaction effects, including fractional quantum Hall states enkner2024b ; bacciconi2023 or correlated states in twisted bilayer graphene Cao in a cavity. While our calculations are restricted to the linear-response regime and the RPA, the description of HPP cavities via the scalar Green’s function provides a natural pathway for extending the framework to a more general diagrammatic formulation that can incorporate non-linear many-body effects. Indeed, our microscopic theory of sub-wavelength cavities provides a rigorous framework to explore how cavity environments can enhance Ebbesen19 ; Sentef_ScienceAdvances_2018 or induce schlawin2019 superconductivity. While previous studies have explored these effects in Fabry-Pérot cavities or within phenomenological models of sub-wavelength cavities, revisiting these results through our fully microscopic approach could provide a more rigorous foundation for superconductivity in cavity-embedded materials. To achieve this, our theory could be interfaced with Eliashberg theory, allowing for a systematic treatment of the strong non-retarded electron-electron interactions induced by the cavity.
@@ VI.1 Acknowledgments
G.M.A acknowledges useful discussion with I. Torre, F.M.D Pellegrino, L. Orsini, J. Sueiro, G. Mazza, Z. Bacciconi, D. De Bernardis, G. Chiriacò, M. Michael, A. Rubio, M. Ruggenthaler and F. Schlawin. G.M.A acknowledge funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie (Grant Agreement No. 101146870 – COMPASS). G.M.A and M.S. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002955 – CONQUER). B.T. acknowledges funding from “la Caixa” Foundation (ID 100010434). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). R.R. and M.P. were supported by the MUR - Italian Ministry of University and Research under the “Research projects of relevant national interest - PRIN 2020” - Project No. 2020JLZ52N (“Light-matter interactions and the collective behavior of quantum 2D materials, q-LIMA”) and by the European Union under grant agreement No. 101131579 - Exqiral and No. 873028 - Hydrotronics. F.H.L.K. acknowledges support from the ERC TOPONANOP under grant agreement no. 726001, the Gordon and Betty Moore Foundation through Grant GBMF12212, and the Government of Spain (FIS2016-81044; PID2019-106875GB-100; Severo Ochoa CEX2019-000910-S [MCIN/ AEI/10.13039/501100011033], PCI2021-122020-2A, and PDC2022-133844-100 funded by MCIN/AEI/10.13039/501100011033). This work was also supported by the European Union NextGenerationEU/PRTR (PRTR-C17.I1) and EXQIRAL 101131579, Fundació Cellex, Fundació Mir-Puig, and Generalitat de Catalunya (CERCA, AGAUR, 2021 SGR 014431656). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Additionally, the research leading to these results received funding from the European Union’s Horizon 2020 under grant agreements no. 881603 (Graphene flagship Core3) and 820378 (Quantum flagship). This material is based upon work supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7047. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force.
@@ Appendix B The density–density response function for graphene in a magnetic field
Employing Eqs. (105,107) we can express the element matrix of the density operator $\hat{n}_{\bm{q}_{\parallel}}$ as: